You purchased 100 shares of General Motors stock at a price of $93.60 one year ago. You sold all stocks today for $99.59. During the year, the stock paid dividends of $3.09 per share. What is your holding period return?

Round the answers to two decimal places in percentage form.
Answer #1

Beginning value =100 * 93.6 = 9,360

Dividends = 100 * 3.09 = 309

Ending value = 100 * 99.59 = 9,959

Holding period return = [(Ending value + dividends - beginning value) / beginning value] * 100

Holding period return = [(9,959 + 309 - 9,360) / 9,360] * 100

Holding period return = 9.70%